OpenStudy (babyslapmafro):

Please show me how to simplify the following expression. (2^n)-2^(n+1) Thanks...

OpenStudy (babyslapmafro):



OpenStudy (anonymous):

\[2^{n}-2^{n+1}=2^n-2\times 2^n\] is a start

OpenStudy (anonymous):

then just like \(x-2x=-x\) you have \(2^n-2\times 2^n=-2^n\)

OpenStudy (babyslapmafro):

for some reason my equation creator is not working and I also can not read what you just wrote.... it looks like this to me \[2^{n}-2^{n+1}=2^n-2\times 2^n\]

OpenStudy (anonymous):

i can read that perfectly what do you see?

OpenStudy (babyslapmafro):

then just like \(x-2x=-x\) you have \(2^n-2\times 2^n=-2^n\)

OpenStudy (anonymous):

try refreshing your browser, maybe you can see it i see it, even when you wrote it

OpenStudy (babyslapmafro):

weird lol...ok ill restart the browser

OpenStudy (anonymous):

just refresh, that might work